### v3

# Tracking the variation of entanglement Rényi negativity : A quantum Monte Carlo study

Abstract
Entanglement entropy has been a powerful tool for analyzing phases and criticality in pure ground states via quantum Monte Carlo (QMC). However, mixed-state entanglement, relevant to systems with dissipation, finite temperature, and disjoint regions, remains less explored due to the lack of efficient numerical methods. In this work, we present a practical and easy-to-implement QMC method within the reweight-annealing framework, enabling efficient computation of the entanglement Rényi negativity (RN) by tracking its variation along given parameter paths. This method is scalable, parallelizable, and well-suited for high-dimensional and large-scale simulations. Applying it to diverse scenarios—including 1D and 2D systems, ground and thermal states, and bipartite and tripartite partitions, not only the information of the underlying conformal field theory is achieved, but the role of entanglement in quantum and thermal phase transitions is revealed.
Introduction.- The increasing connection between condensed matter physics and quantum information science have significantly deepened our understanding of quantum many-body behaviors. Beyond the conventional Landau-Ginzburg-Wilson framework of symmetry breaking and phase transitions Sachdev (1999); Lifshitz and Pitaevskii (2013), quantum entanglement and other intricate microcosmic degrees of freedom have become central in explaining various exotic emergent phenomena Amico et al. (2008); Laflorencie (2016); Zeng et al. (2019). In a bipartite composite system, the entanglement of a pure ground state can be quantified by entanglement entropy (EE), which captures universal information of the system including critical behaviors Kallin et al. (2014); Helmes and Wessel (2014); Zhao et al. (2022a, b), continuous symmetry breaking Metlitski and Grover (2011); D’Emidio (2020); Deng et al. (2023); Wang et al. (2024a), the underlying conformal field theory (CFT) Calabrese and Lefevre (2008); Fradkin and Moore (2006); Casini and Huerta (2007); Ji and Wen (2019); Tang and Zhu (2020) and topological order Kitaev and Preskill (2006); Levin and Wen (2006); Isakov et al. (2011); Nussinov and Ortiz (2009a, b). However, EE fails as a measure of entanglement in mixed states, which frequently arise in many-body physics, such as in finite-temperature Gibbs states, systems with disconnected partitions, and open quantum systems Lifshitz and Pitaevskii (2013); Laflorencie (2016); Yan et al. (2018).
The entanglement logarithmic negativity, referred to as negativity in this letter, is an important entanglement monotone for mixed states that can be computed without requiring optimization procedures Vidal and Werner (2002); Plenio (2005); Audenaert et al. (2002); Eisler and Zimborás (2015); Nobili et al. (2016); Bianchini and Castro-Alvaredo (2016); Eisler and Zimborás (2016); Shapourian et al. (2017); Shapourian and Ryu (2019); Sherman et al. (2016a). Given a bipartite composite system $A\cup B$ in state $\rho$, the negativity that quantifies the entanglement between $A$ and $B$ is defined as $\mathcal{E}:=\ln||\rho^{T_{B}}||$, where $T^{B}$ denotes the partial transpose operating on subsystem $B$, and |||||$\cdot$||| | ⋅ | | is the trace norm, which sums over all absolute values of the eigenvalues of $\rho^{T_{B}}$. According to the positive partial transpose (PPT) criterion Peres (1996); Horodecki (1997); Simon (2000), a nonzero value of negativity indicates the existence of entanglement. While the converse does not always hold, building on the PPT criterion, negativity has been demonstrated as a powerful tool for characterizing mixed-state entanglement Calabrese et al. (2012, 2013a, 2013b); Kulaxizi et al. (2014); Calabrese et al. (2014); De Nobili et al. (2015); Wichterich et al. (2009); Ruggiero et al. (2016); Javanmard et al. (2018); Lee and Vidal (2013); Castelnovo (2013); Wen et al. (2016a, b); Hart and Castelnovo (2018); Lu and Grover (2020a); Sherman et al. (2016b); Lu et al. (2020); Lu (2024). However, for general many-body systems, it is highly challenging to accurately determining the full spectrum of $\rho^{T_{B}}$ to compute negativity. Instead, the moments $\text{tr}[(\rho^{T_{B}})^{n}]$, where $n\geq 1$ is an integer, are more tractable Elben et al. (2020); Calabrese et al. (2012, 2013a, 2013b); Chung et al. (2014a); Alba (2013); Neven et al. (2021). Particularly, the Rényi negativity (RN), a Rényi version of negativity, has emerged as an important quantity for studying mixed-state entanglement Wu et al. (2020a); Wybo et al. (2020); Wang and Xu (2025); Fan et al. (2024); Alba (2013). It is defined as
$R_{n}(\rho):=-\ln\frac{\text{tr}[(\rho^{T_{B}})^{n}]}{\text{tr}(\rho^{n})}$ (1)
Although RN is not an entanglement monotone as it may increase under local operations and classical communications Wybo et al. (2020), it can reflect many important properties of entanglement just as the negativity. For instance, it obeys the area law and the area-law coefficient is singular at the finite temperature critical point Lu and Grover (2019, 2020b); Wu et al. (2020a); Wang and Xu (2025). From the perspective of numerical computation, RN and negativity share a relationship analogous to that between Rényi EE and von Neumann EE, with the former being more tractable.
In recent years, some numerical approaches have been developed to calculate RN or the moments Alba (2013); Shim et al. (2018); Chung et al. (2014a); Wu et al. (2020a); Wang and Xu (2025). However, particularly for large-scale interacting spin/boson systems, there are few quantum Monte Carlo (QMC) works Chung et al. (2014b); Wu et al. (2020b), mainly because the incremental trick commonly used in calculating Rényi EE Hastings et al. (2010); Humeniuk and Roscilde (2012); D’Emidio (2020) becomes much more complex and difficult in the case of RN, since the lowest order of nontrivial RN is third and the boundary condition of imaginary time needs to be incrementally changed from a glued space-time manifold to a twistedly glued one. Therefore, developing an efficient QMC algorithm that is both easy to implement and capable of accurately computing RN is a crucial yet challenging task.
Furthermore, the absence of such algorithms has left many important questions about mixed-state entanglement unexplored. For example, one might naturally assume that the infinite quantum correlation length at a quantum critical point necessarily implies long-range entanglement. However, this is not always the case. Quantum correlation length does not directly measure entanglement: for instance, gapped topological phases can exhibit long-range entanglement despite having finite correlation lengths. Meanwhile, the classical Ising model has no entanglement even at the thermal critical point with infinite correlation length. Nevertheless, the thermal phase transition of a quantum system has both thermal and quantum fluctuations. Due to the existence of quantum fluctuation, the $T_{c}$ of thermal phase transition of a quantum model is usually lower than a pure classical model. It can thus be seen that quantum fluctuations are involved in the thermal phase transition of quantum systems. A fundamental question is what role the entanglement plays in this case.
In this letter, we have filled the technical gap by presenting an efficient QMC algorithm within the reweight-annealing frame Ding et al. (2024); Wang et al. (2024b); Jiang et al. (2024); Ding et al. (2025) and explored several fundamental questions about the entanglement in mixed states.
Figure 1: The horizontal axis represents spatial spins (depicted as circles), while the vertical axis corresponds to imaginary time. Spins of the same color are connected (or glued) along the imaginary time direction by straight lines of the same color. (a) Manifold for $Z_{3}$: The three replicas are sequentially connected along the imaginary time direction for the system $A\cup B$. The environment, denoted by $\overline{A\cup B}$, is traced out by imposing periodic boundary conditions on each replica. (b) manifold for $Z_{3}^{T_{B}}$: The connections among three replicas in subsystem $A$ (the blue part) and the environment $\overline{A\cup B}$ (the yellow part) remain identical to that in the left manifold. However, in subsystem $B$, the connections are not in sequence due to the operation $T_{B}$.
Basic idea.- For convenience, we define two generalized partition functions: $Z_{n}:=\text{tr}(\rho^{n})$ and $Z_{n}^{T_{B}}:=\text{tr}[(\rho^{T_{B}})^{n}]$, which correspond to two types of manifolds (illustrated by $n=3$ in Fig. 1) in the path integral formulation. Thus ${R}_{n}=-\ln Z_{n}^{T_{B}}/Z_{n}$. Here, we define $\rho:=e^{-\beta H}$, which is not normalized in the context of QMC. Typically, $Z_{n}^{T_{B}}$, $Z_{n}$ and ${R}_{n}$ are functions of some real parameter $\lambda>0$ (such as temperature or a Hamiltonian parameter). Our starting point is to consider the difference of $R_{n}$ at two different parameter points $\lambda^{\prime}$ and $\lambda^{\prime\prime}$ ($\lambda^{\prime}<\lambda^{\prime\prime}$), such that
$\begin{split}{R}_{n}(\lambda^{\prime\prime})-{R}_{n}(\lambda^{\prime})=&\ln\frac{Z_{n}^{T_{B}}(\lambda^{\prime})}{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}-\ln\frac{Z_{n}(\lambda^{\prime})}{Z_{n}(\lambda^{\prime\prime})}.\end{split}$ (2)
The ratio $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ in Eq. (2) can be further interpreted as a reweighting operation Ferrenberg and Swendsen (1988, 1989); Troyer et al. (2004); Dai and Xu (2025). If express $Z_{n}^{T_{B}}(\lambda)=\sum_{i}W_{i}(\lambda)$, where $\{W_{i}(\lambda)\}$ are the weights, then the ratio can be viewed as the average of weight ratios, since $Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})=\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$, where $W(\lambda)\in\{W_{i}(\lambda)\}$ and the expectation is computed with respect to samples from $Z_{n}^{T_{B}}(\lambda^{\prime\prime})$. The specific form of the estimator $\langle W(\lambda^{\prime})/W(\lambda^{\prime\prime})\rangle_{Z_{n}^{T_{B}}(\lambda^{\prime\prime})}$ depends on the QMC method we use, which transforms the quantum degrees of freedom in $Z_{n}^{T_{B}}(\lambda^{\prime})$ to some classical ones. For illustration, we use the stochastic series expansion (SSE) method in this letter Sandvik (1998, 2003); Yan et al. (2019); Yan (2022), and the corresponding estimators can be found in the Supplemental Material sup (see also references Sandvik (1998, 2003); Yan et al. (2019); Yan (2022); Zhou et al. (2022, 2023); Yan et al. (2023); Calabrese et al. (2013a); Laflorencie (2016) therein). Similarly, we can apply the same reasoning to $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$. It is important to emphasize that the view of reweighting is crucial, as it allows us to estimate the difference $[{R}_{n}(\lambda^{\prime\prime})-{R}_{n}(\lambda^{\prime})]$ from Eq. (2) by estimating $[Z_{n}^{T_{B}}(\lambda^{\prime})/Z^{T_{B}}_{n}(\lambda^{\prime\prime})]$ and $[Z_{n}(\lambda^{\prime})/Z_{n}(\lambda^{\prime\prime})]$ in QMC simulations. Additionally, to avoid overflow/underflow issues, we have to store their logarithms in computers.
The procedure described above is sufficient to derive the relative values of $R_{n}(\lambda)$ at different $\lambda^{\prime\prime}$ when $\lambda^{\prime}$ is fixed at $\lambda_{0}$, called a reference point. This allows for the study of various behaviors, such as locating extreme values and singularities of $R_{n}(\lambda)$. Moreover, if ${R}_{n}(\lambda_{0})$ is known, we can compute the exact values of ${R}_{n}(\lambda)$. Typically, finding such a reference point is straightforward. For example, when $\lambda=\beta\equiv 1/T$, the inverse temperature, we can set $\beta_{0}=0$, which corresponds to a completely disordered system where both classical and quantum correlations are destroyed, giving ${R}_{n}(\beta_{0})=0$. In practice, we use a sufficiently small value (e.g. $\beta_{0}=10^{-8}$) in our simulations, and the adequacy of the choice can be verified by considering a smaller one to check the convergence. Similarly, when $\lambda$ represents the coupling strength of interacting terms in a Hamiltonian related to the entangled boundary, setting $\lambda_{0}=0$ would make the two subsystems completely independent, resulting in ${R}_{n}(\lambda_{0})=0$ again. Therefore, depending on our choice of $\lambda$, both finite and zero temperature properties of $R_{n}$ can be explored. For illustrative purpose, we now set $\lambda=\beta$, and other choices of $\lambda$ can be similarly discussed. By identifying $\lambda_{0}\equiv\beta_{0}\equiv 0$, Eq. (2) can be reformulated as ${R}_{n}(\beta)=\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}-\ln\frac{Z_{n}(0)}{Z_{n}(\beta)}$.
Importance sampling and annealing scheme.- Although the ideas of reweighting and reference point are simple, simulations become inefficient if $\beta$ is far away from $\beta_{0}$ in the parameter space, due to the significant differences between the distributions $Z_{n}^{T_{B}}(\beta)$ and $Z_{n}^{T_{B}}(\beta_{0})$, requiring an excessive number of samples. To address this, we divide $[\beta_{0},\beta]$ into $m$ small subintervals, i.e. $\ln\frac{Z_{n}^{T_{B}}(0)}{Z_{n}^{T_{B}}(\beta)}=\sum_{k=1}^{m}\ln\frac{Z_{n}^{T_{B}}(\beta_{k-1})}{Z^{T_{B}}_{n}(\beta_{k})}$ and $\ln\frac{Z_{n}(0)}{Z_{n}(\beta)}=\sum_{k=1}^{m}\ln\frac{Z_{n}(\beta_{k-1})}{Z_{n}(\beta_{k})}$, where $\beta_{m}\equiv\beta$. Then, efficient estimation is ensured when adjacent $\beta_{k}$ values are close to satisfy the importance sampling Neal (2001). In addition, the method’s effectiveness depends on how $[\beta_{0},\beta]$ is divided, as the computational cost also scales with the number of subintervals. Simple approaches, such as equal divisions or geometric progressions with large/small ratios, are inefficient since the required number of subinterval density varies across parameter regions and different system sizes.
To address the problem, we adopt the annealing scheme, which requires a polynomial number of subintervals and has demonstrated excellent performance in practical partition function calculations Ding et al. (2024). The core idea is to approximately restrict $[Z_{n}^{T_{B}}(\beta_{k-1})/Z^{T_{B}}_{n}(\beta_{k})]$ and $[Z_{n}(\beta_{k-1})/Z_{n}(\beta_{k})]$ to be some constant $\epsilon\lesssim 1$ for all sizes and $\beta_{k}$, to ensure the importance sampling. With some calculations, we find that the total number of subintervals $F(L)$ required scales as $F(L)\approx(\beta-\beta_{0})\Lambda nL^{d}/|\ln\epsilon|$, where $d$ is the system dimension, $L$ is the length of system, and $\Lambda$ is a constant related to the energy density. Notably, an incorrect choice of $\Lambda$ just effectively changes the value of $\epsilon$, making $\epsilon$ the sole hyperparameter in the algorithm. Similarly, one can discuss when $\lambda$ is some other parameter in the Hamiltonian, and the division number also scales polynomially with the system size. Further details, including the derivation of $F(L)$ can be found in the Supplemental Material sup .
Ground states of tripartite 1D antiferromagnetic dimerzied Heisenberg model (ADHM).-
Figure 2: With $\beta=L$ to extrapolate to the ground state: (a) The variation of $R_{3}(K)$ with respect to the parameter $K$ for the tripartite (1+1)D ADHM ring; (b) The central charge $c=1.02(2)$ of the underlying CFT for the Heisenberg model is fitted; (c) The variation of $R_{3}(h)$ with respect to the parameter $h$ for the (1+1)D transverse field Ising model (TFIM) under half bipartition; (d) The central charge of the underlying CFT at the critical point is fitted to be $c=0.51(3)$. Notably, the data of $L=16$, which exhibits significant finite-size effects and clearly deviates from the trend, is excluded from the fitting.
For the first example, we consider a 1D ring of ADHM whose Hamiltonian is $H_{\text{ADHM}}=\sum_{i\in\text{odd}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+K\sum_{i\in\text{even}}\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}$. We partition the system into three connected and equally sized regions: subsystems $A$, $B$, and the environment $\overline{A\cup B}$, with boundary bonds having a coupling strength of $K$. When $K=0$, the three regions $A$, $B$, and $\overline{A\cup B}$ are completely independent, thus we take $K_{0}=0$ as the reference point and gradually increase $K$ from $K_{0}$ to $K=1$ in simulations. By tracing out the environment $\overline{A\cup B}$ according to Fig. 1, the entanglement between subsystem $A$ and $B$ is in the context of mixed state. As shown in Fig. 2(a), the entanglement quantified by $R_{3}$ increases monotonically when increasing $K$. This indicates that the entanglement between $A$ and $B$ arises from the Heisenberg interaction at the boundaries. Moreover, at $K=1$, the model is described by a $(1+1)$D CFT, and $R_{n}$ satisfies Calabrese et al. (2012, 2013a)
$R_{n}=\begin{cases}\frac{c}{12}(n-\frac{1}{n})\ln\bigg{(}\frac{L}{\pi a}\frac{\sin{\frac{\pi l_{1}}{L}}\sin{\frac{\pi l_{2}}{L}}}{\sin{\frac{\pi(l_{1}+l_{2})}{L}}}\bigg{)}+\mathcal{O}(1),n\in\text{odd}\\ \frac{c}{6}(\frac{n}{2}-\frac{2}{n})\ln\bigg{(}\frac{L}{\pi a}\frac{\sin{\frac{\pi l_{1}}{L}}\sin{\frac{\pi l_{2}}{L}}}{\sin{\frac{\pi(l_{1}+l_{2})}{L}}}\bigg{)}+\mathcal{O}(1),n\in\text{even}\end{cases}$ (3)
where $c$ is the central charge, $a$ is the lattice constant, $L$ is total length of the ring, and $l_{1}$ and $l_{2}$ are the lengths of subsystems $A$ and $B$, respectively. Note that the standard negativity $\mathcal{E}$ can be obtained through taking $n\to 1$ for even $n$. For $n=3$, we therefore have $R_{3}\propto(2c/9)\ln L$. As shown in Fig. 2(b), by fitting $R_{3}(K=1)$ with the data, we successfully extract $c=1.02(2)$, which is consistent with the theoretical value $1$ Blöte et al. (1986).
Ground states of bipartite 1D TFIM.- We next consider the TFIM, and the Hamiltonian is given by $H=-\sum_{\langle ij\rangle}Z_{i}Z_{j}-h\sum_{i}X_{i}$, where $\langle ij\rangle$ denotes the nearest neighbor. For the 1D ring, we consider half bipartition without the environment. By fixing $J=1$, the quantum critical point (QCP) is at $h_{c}=1$, which is also described by a CFT with central charge $c_{\text{Ising}}=1/2$. We take $h_{0}=0$ to be the reference point, and Fig. 2(c) shows that $R_{3}(h)$ is maximized at the QCP. When increasing the system length $L$, it is observed that $R_{3}(L,h_{c})$ diverges. In addition, in the case of half bipartition, $R_{n}$ satisfies
$R_{n}(h_{c})=\begin{cases}\frac{c}{6}(n-\frac{1}{n})\ln\bigg{(}\frac{L}{\pi a}\sin\frac{\pi l}{L}\bigg{)}+\mathcal{O}(1),&n\in\text{odd}\\ \frac{c}{3}(\frac{n}{2}-\frac{2}{n})\ln\bigg{(}\frac{L}{\pi a}\sin\frac{\pi l}{L}\bigg{)}+\mathcal{O}(1),&n\in\text{even}\end{cases}$ (4)
for a (1+1)D CFT, where $l$ is the length of subsystems $A$ and $B$. With Eq. (4), the central charge is obtained to be $c=0.51(3)$, which also accords with the theoretical value. In both the tripartite 1D ADHM and the bipartite 1D TFIM, the logarithmic corrections to $R_{n}$ at the CFT points deviate from the expected relation $R(2l)=2R(l)$, signaling the presence of long-range entanglement Lu et al. (2020); Lu and Grover (2020b); Lu (2024). In contrast, a system with only short-range entanglement would yield an extensive $R_{n}$, with no subleading terms beyond the area law.
Figure 3: For the 2D TFIM on the square lattice: (a) fixing $h=3.04438$, the variation of $R_{3}(J)$ as a function of $J$, where $\beta=L$ to extrapolate to the ground state; (b) the area law correction $-\gamma(J)$ of $R_{3}(J)$ for $J\in[0,1.2]$. The inset panel shows the fitting of area law with the maximum value of $R_{3}(L,J)$ for each $L$, and the area law coefficient is $\gamma=-0.225(49)$; (c) fixing $J=1$ and $h=2.75$, the variation of $R_{3}(\beta)$ as a function of $\beta$. (d) the area law correction $-\gamma(\beta)$ of $R_{3}(\beta)$ for $\beta\in(0,0.4]$, and the inset shows the fitting of the area law at $\beta=\beta_{c}/3$, where $-\gamma(\beta_{c})=0.002(7)$;
Ground states and Gibbs states of bipartite 2D TFIM.- We similarly investigated $R_{3}$ of the 2D TFIM on a $L\times L$ square lattice with periodic boundary condition, considering the cornerless half-bipartition. The QCP of its ground state occurs at $h/J=3.04438(2)$ Blöte and Deng (2002). For convenience, we fix $h=h_{c}\equiv 3.04438$ and tune $J$, so that the QCP is positioned at $J_{c}\approx 1$. To study the ground-state entanglement, we consider $R_{3}$ as a function of $J$ and take $J_{0}=0$ as the reference point. Furthermore, when $h/J<3.04438$, increasing the temperature can induce a thermal phase transition from a ferromagnetic to a paramagnetic phase Hesselmann and Wessel (2016). Thus we can set a large $h$ to highlight the quantum fluctuations, and study the finite-temperature mixed-state entanglement across a thermal critical point (TCP). We choose $J=1$ and $h=2.75$, and the TCP located at $\beta_{c}=1/T_{c}=1.0874(1)$ Wu et al. (2020a).
For the ground states, Fig. 3(a) illustrates the behavior of $R_{3}(J)$ as a function of $J$ for different system sizes. Similar to the 1D TFIM, $R_{3}(J)$ reaches a maximum and diverges at the QCP, indicating that the QCP exhibits stronger quantum correlations compared to the gapped phases. In the gapped phases, no long-range entanglement is expected, and we should have $R_{3}\sim aL-\gamma$ with $\gamma=0$. Near the critical point, we observe that the area law breaks down, with $R_{3}$ for small system sizes exceeding that of larger sizes, as shown in Fig. 3(a). However, as the system moves further from the critical point, the area law is restored. If we forcibly fit $R_{3}$ to the area law, the constant term $\gamma$ fluctuates significantly near the critical point, as shown in Fig. 3(b). In other regions, it remains zero within error bars. In principle, $R_{n}$ can be understood as the difference of free energies between two space-time manifolds with different boundary conditions. Since the only distinction lies at the boundary, it should follow an area law. The deviation from the area law in our numerical data is due to finite-size effects in both space and time, as the gaplessness related to the phase transition requires larger system sizes. Further details on finite-size effects are provided in the Supplemental Material sup .
To reduce finite-size effects, we use the peaks of $R_{3}(J)$ across different system sizes to fit the area law at the QCP, as shown in the inset of Fig. 3(b) [^1] For large sizes, this result converges to that obtained by fitting at a fixed parameter point.. Its constant correction extracted from the fit is $\gamma=-0.225(49)$  [^2] Notice we use $R_{3}\sim aL-\gamma$ to fit, thus the final constant correction implies a positive sub-leading correction., which signifies a breakdown of the local behavior of RN and reflects the long-range entanglement at the QCP here. This constant correction is consistent with the theoretical prediction Metlitski et al. (2009). In addition, RN inherently removes the contribution of ground state degeneracy due to the denominator in its definition, whereas EE captures it in the correction term as $\gamma=-\ln(N_{\text{deg}})$, where $N_{\text{deg}}$ denotes the ground state degeneracy Stéphan et al. (2009). In this example of ferromagnetic phase, we have $\gamma=0$ while $N_{\text{deg}}=2$. This reflects the advantage of RN in computing mixed-state properties, where it naturally filters out degeneracy effects.
For the Gibbs states, Fig. 3(c) shows that $R_{3}(\beta)$ increases monotonically from the infinite high temperature $\beta_{0}$ to $\beta=0.4$. Notably, the three replicas effectively scale the inverse temperature by a factor of $1/3$ in partition function $Z_{3}$, and the related critical properties are shifted to $\beta_{c}/3$  [^3] The term $\mathrm{tr}(\rho^{3})\propto\text{tr}(e^{-3\beta H})$ will lead a singularity at $\beta_{c}/3$. This has been confirmed in the calculation of the divergence of the first derivative of $a$, the area-law coefficient of $R_{3}$, in the quantum spherical model and 2D TFIM Wu et al. (2020a). As seen in Fig. 3(c), at the critical point $\beta_{c}/3$, no significant behavior associated with the TCP is observed for $R_{3}$ in our calculations. Furthermore, we find that $R_{3}$ perfectly obeys the area law and $\gamma\sim 0$ for all the $\beta$ we computed, as shown in Fig. 3(d). Despite choosing $h=2.75$, close to the QCP and dominated by strong quantum fluctuations, our findings indicate that thermal phase transitions remain short-range entangled, highlighting a key distinction from the long-range entanglement inherent in quantum criticality. Moreover, the increase in $R_{3}$ when decreasing the temperature reflects an enhancement of quantum properties. However, the quantum correlations remain short range ($\gamma\sim 0$), indicating that the phases beyond or below the TCP are both trivial in terms of quantum correlation.
In addition, we discuss the derivatives of $R_{n}$ with respect to some parameter $\lambda$. Given that $R_{n}(\lambda)=-\ln Z_{n}^{T_{B}}(\lambda)+\ln Z_{n}(\lambda)$, it follows that if the $k$th-order derivatives of both effective free energy terms are non-singular at $\lambda$, then $d^{k}R_{n}(\lambda)/d\lambda^{k}$ must also remain non-singular. Moreover, since the difference between the two systems associated with $Z_{n}^{T_{B}}$ and $Z_{n}$ arises solely from their boundary conditions, their critical points $\lambda_{c}$ should remain identical in the thermodynamic limit. Consequently, any singularity in a higher-order derivative of $R_{n}$ must occur precisely at $\lambda_{c}$.
Summary and discussions.- We have introduced an efficient algorithm for calculating Rényi negativity in quantum Monte Carlo simulations, built on the reweight-annealing framework. This algorithm is easy to implement, has polynomial computational complexity, and is naturally parallelizable. By enabling the calculation of Rényi negativity along a full parameter trajectory in parameter space, our method proves highly versatile. It is not only valuable for investigating entanglement behavior across different phases and near critical points, but also for fitting correction terms in area laws to extract universal quantities such as central charge and topological order in quantum many-body systems. We have validated the approach in various scenarios, including 1D and 2D systems, ground and Gibbs states, as well as bipartition and tripartition setups. Our results of the Rényi negativity further reveals a detailed understanding of the intrinsic mechanisms governing thermal and quantum critical points in many-body systems. An intriguing direction for future research is to investigate the role of entanglement in strong-to-weak spontaneous symmetry breaking in mixed quantum states Lessa et al. (2025a, b).
Acknowledgements.- The authors sincerely thank Tsung-Cheng Lu for kindness and valuable discussions. The authors also thank Fo-Hong Wang, Wei Zhu, Rui-Zhen Huang, and Ying-Jer Kao for their helpful discussions. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. ZW is supported by the China Postdoctoral Science Foundation under Grants No.2024M752898. BBM acknowledge the Natural Science Foundation of Shandong Province, China (Grant No. ZR2024QA194) and NSFC Grant No. 12247101. The authors thank the high-performance computing center of Westlake University and the Beijng PARATERA Tech Co.,Ltd. for providing HPC resources.